OpenStudy (anonymous):

write the slope-intercept equation for the line with the given slope that contains the given point. m= -3; (-2,1)

OpenStudy (anonymous):

help please!!!

OpenStudy (anonymous):

y=-3x+b 1=-3(-2)+b 1=6+b b=-5 y=-3x-5

OpenStudy (anonymous):

thanks